What is the probability your friend will pass the exam. There are 20 true/false items on the exam. Each question has the same weight and 60% or more correct answers results in a passing grade.
1
Expert's answer
2020-03-02T11:28:28-0500

Consider "X"  as the number of questions answered correctly among the 20 questions: "X\sim B(n,p)"


"P(X=x)=\binom{n}{x}p^x(1-p)^{n-x}"

Given that "p=0.5, n=20"

"0.6\cdot20=12"


"P(X=12)=\binom{20}{12}0.5^12(1-0.5)^{20-12}=0.120134"

"P(X=13)=\binom{20}{13}0.5^{13}(1-0.5)^{20-13}=0.073929"

"P(X=14)=\binom{20}{14}0.5^{14}(1-0.5)^{20-14}=0.036964"

"P(X=15)=\binom{20}{15}0.5^{15}(1-0.5)^{20-15}=0.014786"

"P(X=16)=\binom{20}{16}0.5^{16}(1-0.5)^{20-16}=0.004621"

"P(X=17)=\binom{20}{17}0.5^{17}(1-0.5)^{20-17}=0.001087"

"P(X=18)=\binom{20}{18}0.5^{18}(1-0.5)^{20-18}=0.000181"


"P(X=19)=\binom{20}{19}0.5^{19}(1-0.5)^{20-19}=0.000019"

"P(X=20)=\binom{20}{20}0.5^
{20}(1-0.5)^{20-20}=0.000001"

"P(X\geq12)=P(X=12)+P(X=13)+P(X=14)+""+P(X=15)+P(X=16)+P(X=17)+P(X=18)+"

"+P(X=19)+P(X=20)\approx" "0.251722"


The probability my friend will pass the exam is "0.2517" "(\approx25\%)"